OpenStudy (anonymous):

Written in slope-intercept form, the equation x + 2y = 4 is

OpenStudy (anonymous):

y=-1/2x + 2

OpenStudy (anonymous):

do you understand how phanta_seea got that answer ?

OpenStudy (anonymous):

it is the correct answer... x + 2y = 4 (subtract x from both sides) 2y = -x + 4 (now divide by 2) y = -1/2x + 2

OpenStudy (anonymous):

i understood i was just double checking if you would also divide out the 4